# SpecTool: A Benchmark for Characterizing Errors in Tool-Use LLMs

### v1

@@ Abstract
Evaluating the output of Large Language Models (LLMs) is one of the most critical aspects of building a performant compound AI system. Since the output from LLMs propagate to downstream steps, identifying LLM errors is crucial to system performance. A common task for LLMs in AI systems is tool use. While there are several benchmark environments for evaluating LLMs on this task, they typically only give a success rate without any explanation of the failure cases. To solve this problem, we introduce SpecTool, a new benchmark to identify error patterns in LLM output on tool-use tasks. Our benchmark data set comprises of queries from diverse environments that can be used to test for the presence of seven newly characterized error patterns. Using SpecTool , we show that even the most prominent LLMs exhibit these error patterns in their outputs. Researchers can use the analysis and insights from SpecTool to guide their error mitigation strategies.
Machine Learning, ICML
*skokane@salesforce.com
Salesforce AI Research, USA
@@ 1 Introduction
An emerging use case for LLMs in AI systems is tool use. When LLMs are equipped with context and a list of tools, such as APIs, databases, and applications, LLMs can generate a sequence of calls to the available tools to solve a given task. Using LLMs to perform tasks via APIs involves more complex reasoning and instruction following abilities of LLMs, and there are several emerging methods that aim to improve those abilities in order to help LLMs adapt to new tasks and situations. These methods (Zhang et al., 2024; Li et al., 2023) help enhance the model’s understanding of user preferences, along with its skills in logical reasoning, with the goal of increasing their overall effectiveness. The basic technique involves using guiding prompts to provide LLMs with instructions and information about the context, enabling them to generate actions to solve complex tasks. Besides, some techniques (Rafailov et al., 2024; Jing et al., 2019) focus on dedicated training methods to turn LLMs into highly capable agents.
@@
In an effort to address the limitations of the existing benchmarks, we present SpecTool , a benchmark created to characterize common errors in tool-use LLMs and provide detailed diagnostic feedback to help improve them. SpecTool comprises of 10 distinct environment categories, including Tools, Movies, Travel, Sports, Entertainment, Data, Social, Media, Weather and Video Images, making it one of the most comprehensive evaluation environments. It also incorporates over 30 different tasks specifically aimed at tool agents, such as entertainment, sports, and weather tasks.
@@
Furthermore, the SpecTool evaluation dataset comprised of 150 queries can be employed to identify these error patterns in the output of LLMs during tool-use tasks. These queries have been annotated by humans to highlight multiple pathways to reach a given objective. Subsequently, we showcase the capacity of SpecTool to identify error patterns in a case study involving various leading LLMs.
@@
  • We introduce SpecTool , a comprehensive benchmark covering 10 environment categories and over 30 tasks, specifically designed to provide detailed diagnostic feedback on tool-use tasks in LLMs.
@@
  • We provide a 150-query human-annotated dataset to detect error patterns in LLM outputs and demonstrate SpecTool ’s effectiveness through a case study involving several leading LLMs.
Table 1: Comprehensive Comparison of SpecTool across critical evaluation requirements. SpecTool supersedes on all the provided principles, focusing realistic multi-turn interactions.
Points Toolbench Agentboard MintBench Gorilla SpecTool
Multi-Turn Interactions
@@ 2 Related Work
The current benchmarks exhibit notable strengths and weaknesses in evaluating model performance as a tool-use agent. Benchmarks like Gorilla (Patil et al., 2023), AgentBoard (Ma et al., 2024), ToolBench (Guo et al., 2024), ToolEyes (Ye et al., 2024), and MintBench (Wang et al., 2023) are remarkably utilised in examining model interaction ability, reasoning and planning ability. While some benchmarks like ToolBench , AgentBoard , and MintBench are effective in managing multi-turn interactions, others such as Gorilla excel in handling diverse queries, especially when dealing with constraints and irrelevant information. Gorilla also offers detailed error pattern analysis and are well-suited for complex queries that require a wide range of API capabilities but is limited to single-turn interactions only.
However, these evaluations often fail to consider the model’s existing knowledge, leading to unnecessary and repetitive API calls without specific prompt instructions. Furthermore, the benchmarks like AgentBoard place undue emphasis on the sequence of subgoals, sometimes misjudging a model’s progress when redundant subgoals are involved. This rigid focus can result in an inaccurate assessment of the model’s true capabilities in task completion.
We address all these current issues in the existing benchmark, where our focus is more on assessing the quality of the action generated by the model given the constraints in the query. Our emphasis is on understanding some critical errors of the model such as hallucinations in actions w.r.t API names or argument names, performing redundant repetitive calls, invalid format issues. These issues are commonly experienced during training or fine-tuning of an LLM for agentic tasks and it is critically important to quantify such issues to get a deeper insight into the improvements required in the model.
@@ 3 Error Patterns in Tool-Use
It’s widely understood that LLMs are prone to output errors such as hallucinations, inconsistencies, errors, etc (Wei et al., 2022, 2024; Lu et al., 2024). What is less studied are the errors exhibited in tool-use scenarios. In tool-use scenarios, typically the LLM outputs one or more function calls, including necessary arguments for the execution of the functions. For example, an LLM output for using a web search API to learn about the latest fashion can be as the following: $Search[query=``latestfashion",top_{k}=10]$, where $Search$ indicates the tool and $query$ and $top_{k}$ are the context-specific arguments for executing the tool call.
@@
Note that these tool calls are particularly brittle - even a slight change to an argument, a missing argument, or incorrect tool call can produce drastically different and incorrect results downstream. Given the importance of catching these output errors, in this work, we characterize systematic errors generated by LLMs on tool-use tasks into seven critical error types.
@@
Insufficient API Calls (IAC): The LLM is unable to generate sufficient API calls, hence unable to completely fulfill the tasks provided in a query.
@@
Incorrect Argument Value (IAV): The LLM generates incorrect argument values. This also includes exclusion of necessary arguments.
Incorrect Argument Name (IAN): The LLM hallucinates argument names.
Incorrect Argument Type (IAT): The LLM generates incorrect argument type.
@@
Repeated API Calls (RAC): The LLM generates the exact same API Call repeatedly, leading to redundant calls.
Incorrect Function Name (IFN) : The LLM hallucinates function names, that are not included in the API list.
Invalid Format Error (IFE) : The LLM is incapable of following the appropriate format instructions provided for parsing.
Figure 1: Query Generation Workflow for SpecTool.
4 SpecTool Dataset
@@
The brittle nature of LLM-generated tool calls highlights the need to quickly discover errors in LLM-generated outputs. To quickly discover tool-calling errors in LLMs, we propose a new dataset, SpecTool . It is created through a series of effective steps aimed at generating complex queries.
@@
The process begins by gathering seed queries for each API environment, sourced from benchmark tests conducted on toolsets like ToolBench (Guo et al., 2024) and AgentBoard (Ma et al., 2024). These initial seed queries are then augmented using the following methods. See figure 1 for an overview.
@@
Constraint based Query Generation: We employ a deterministic method to discern the options available for each argument within every API call. Utilizing prompt engineering techniques, we introduce detailed descriptions of each argument and its corresponding options to enhance the original query within GPT-4. First, we include one argument along with its option to create a more detailed query. Then, we improve our method by adding multiple arguments and their options, making the query even more complex.
@@
Relevance Data based Query Generation: On seed queries, we instruct GPT-4 to introduce additional, unnecessary information around the keywords while keeping the context of the query unchanged.
We collected around 150 augmented queries and annotated the expected API calls. Evaluating these queries across different models will help us better understand the error patterns generated by LLMs.
The table 2 describes the distribution of queries across different environment APIs. We also list down the average number of interactions alongwith the queries sampled per environment after augmentation.
Table 2: Distribution of SpecTool Dataset. Sampled Queries indicates the queries generated using the Workflow. Avg. Interactions indicates the average steps executed to solve the query. Avg. APIs indicates the average no. of API options for respective environment.
Environments Sampled Queries Avg. Interactions Avg. APIs
@@
SpecTool 150 6.8 6.2
Query Error Model Output
I need jokes sorted by score in ascending order of the category food? Can you provide me with that? IAV MLlama GT: getJokeCategory(category=‘food’, sortby= ‘asc’) MO: getJokeCategory(category=‘workplace’, sortby = ‘asc’)
Can you provide me with the genre, runtime, IMDb rating, language for the movie ‘Endgame’? Also, please include the streaming platforms available in the US. IAN GPT-4 GT: getMovieAdvance(genre = ‘horror’, startyear = 1970, endyear = 2020, minimdb = 7) MO: getMovieAdvance(genre=‘horror’, releasestart=1970 , releaseend = 2020, imdbrating = 7)
I need the titles and their release years for horror movies with a minimum IMDb rating of 7 that were released between 1970 and 2020. IAC GPT-3.5 GT: getMovie(title = ‘Endgame’, type= ‘movie’),            getMovieProvider(region = ‘US’) MO: getMovie(title = ‘Endgame’, type=‘movie’)
I need to check if the domain ‘business.com’ is available or not IAT xLAM GT: checkDomainAvailability(domain = ‘business.com’, availableonly=True) MO: checkDomainAvailability(domain= ‘business.com’ , availableonly=”True”)
Figure 2: A systemic workflow of an agent interacting with the environment based on prior information and the goal.
5 Method
@@
As shown in figure 2, our benchmark follows a deterministic framework where the agent will begin with Instruction i, Query q, and a sequence of api-list (resembling to the action space) $A$ provided to successfully solve the provided query. The agent then based on the policy of the language model will perform an action $a_{t}$ in the current $s_{t}$, leading to a determined state transition $s_{t+1}$.
A constructive feedback mechanism f is also incorporated to help assist the agent for the following issues:
  • Verify whether the generated action is correctly parsed and adheres to the format instructions provided in the prompt.
  • Determine if the generated action is included in the specified action space. If it is not, enumerate the list of available actions that can be used to resolve the query.
  • Assess whether the generated action is invoked with the appropriate arguments. If the arguments are incorrect, provide a detailed list of the available arguments for the chosen action, along with their respective descriptions.
  • Ensure that the generated action is called with arguments of the correct type. If any discrepancies are found, specify the correct argument type required for proper execution of the function.
@@ 6 Metrics
Current benchmarks tend to emphasize the final outcome of models, often overlooking the errors that occur during intermediate steps. Benchmarks such as AgentBoard and MintBench attempt to address this by imposing constraints on the expected outcomes at each step and the sequence in which goals should be achieved. For instance, AgentBoard defines subgoals for each query set in a prescribed order, with these subgoals representing the expected API responses based on the agent’s actions. However, these benchmarks do not consider that the model might possess prior information allowing it to skip certain subgoals, potentially leading to an inaccurate assessment of the model’s performance due to its deviation from the prescribed subgoal sequence. Additionally, these benchmarks lack detailed analysis of specific errors in the action execution, thus failing to provide comprehensive insights into the minor mistakes made by the model.
In this study, we concentrate our efforts on offering deeper insights into specific model inefficiencies. We do this by providing critical metrics related to API Call, thus fostering a comprehensive understanding of the necessary improvements in the models. At present, we have accurately verified and labeled a compilation of 150 queries. These queries are unique, stemming from diverse environments as shown in table 2, and each boasting multiple expected API Call trajectories. We employ these queries to subsequently compute metrics for a variety of models, thereby identifying their respective error patterns. Each metric is defined as follows: N represents the total number of maximum steps the model is permitted to effectively process the query, while GT denotes the Ground Truth labels for each individual query. Our calculation of these metrics strictly follows the sequence outlined above.
@@
  • For error patterns IFN, IAN, IAT, IFE and RAC we calculate each error’s accuracy as: $\frac{N-\mathrm{N_{e}}}{N}$, where $N_{e}$ is the number of API calls executing the error pattern.
  • For error patterns IAC and IAV, we use the labeled ground truths of each query and calculate these metrics across each labeled trajectory: $\min_{g\in GT}\left(\frac{N-N_{g}}{N}\right)$, where $N_{g}$ is the number of API calls with errors w.r.t. the ground truth g.
@@
It is crucial to emphasize that while the metrics indicate error patterns, their calculations are designed to be directly proportional to the success rate. Higher scores on the error metrics represent fewer issues generated by a model in that respective error pattern. We calculate the success rate of each query by comparing its final output with the ground truth. If a quantitative ground truth is unavailable, success is determined by the model’s ability to conclude using the finish API with a correct response. Error metrics are then incorporated with appropriate weightage in both cases.
Table 4: Overall Results of SpecTool Benchmark across several performant LLMs on the proposed error metrics. IAC, IAV, RAC, IFE, IAT, IAN, IFN are error metrics ocurring in each API Calls.

### v2

# ToolScan: A Benchmark for Characterizing Errors in Tool-Use LLMs

@@ Abstract
Evaluating Large Language Models (LLMs) is one of the most critical aspects of building a performant compound AI system. Since the output from LLMs propagate to downstream steps, identifying LLM errors is crucial to system performance. A common task for LLMs in AI systems is tool use. While there are several benchmark environments for evaluating LLMs on this task, they typically only give a success rate without any explanation of the failure cases. To solve this problem, we introduce ToolScan, a new benchmark to identify error patterns in LLM output on tool-use tasks. Our benchmark data set comprises of queries from diverse environments that can be used to test for the presence of seven newly characterized error patterns. Using ToolScan, we show that even the most prominent LLMs exhibit these error patterns in their outputs. Researchers can use these insights from ToolScan to guide their error mitigation strategies.
@@ 1 Introduction
An emerging use case for LLMs in AI systems is tool use. When LLMs are equipped with context and a list of tools, such as APIs, databases, and applications, LLMs can generate a sequence of function calls to solve a given task. Using LLMs to perform tasks via APIs involves more complex reasoning and instruction following abilities of LLMs, and there are several emerging methods that aim to improve those abilities in order to help LLMs adapt to new tasks and situations. These methods (Zhang et al., 2024; Li et al., 2023) help enhance the model’s understanding of user preferences, along with its skills in logical reasoning, with the goal of increasing their overall effectiveness. The basic technique involves using guiding prompts to provide LLMs with instructions and information about the context, enabling them to generate actions to solve complex tasks. Besides, some techniques (Rafailov et al., 2024; Jing et al., 2019) focus on dedicated training methods to turn LLMs into highly capable agents.
@@
In an effort to address the limitations of the existing benchmarks, we present ToolScan, a benchmark created to characterize common errors in tool-use LLMs and provide detailed diagnostic feedback to help improve them. ToolScan comprises of 10 distinct environment categories, including Tools, Movies, Travel, Sports, Entertainment, Data, Social, Media, Weather and Video Images, making it one of the most comprehensive evaluation environments. It also incorporates over 30 different tasks specifically aimed at tool agents, such as entertainment, sports, and weather tasks.
@@
Furthermore, the ToolScan evaluation dataset comprised of 150 queries can be employed to identify these error patterns in the output of LLMs during tool-use tasks. These queries have been annotated by humans to highlight multiple pathways to reach a given objective. Subsequently, we showcase the capacity of ToolScan to identify error patterns in various leading LLMs.
@@
  • We introduce ToolScan, a comprehensive benchmark covering 10 environment categories and over 30 tasks, specifically designed to provide detailed diagnostic feedback on tool-use tasks in LLMs.
@@
  • We provide a 150-query human-annotated dataset to detect error patterns and demonstrate ToolScan’s effectiveness through a case study involving several leading LLMs. Number of tool-use queries in our dataset are similar to other datasets while providing better coverage on diversity, feedback mechanism etc.
Table 1: Comprehensive Comparison of ToolScan across critical evaluation requirements. ToolScan supersedes on all the provided principles, focusing on realistic multi-turn interactions.
Points ToolBench AgentBoard MintBench Gorilla ToolScan
@@ 2 Related Work
The current benchmarks exhibit notable strengths and weaknesses in evaluating model performance as a tool-use agent. Benchmarks like Gorilla (Patil et al., 2023), AgentBoard (Ma et al., 2024), ToolBench (Guo et al., 2024), ToolEyes (Ye et al., 2024), and MintBench (Wang et al., 2023) are remarkably utilized in examining model interaction ability, reasoning and planning ability. While some benchmarks like ToolBench, AgentBoard, and MintBench are effective in managing multi-turn interactions, others such as Gorilla excel in handling diverse queries, especially when dealing with constraints and irrelevant information. Gorilla also offers some error pattern analysis and are well-suited for complex queries that require a wide range of API capabilities but has limited diverse tasks.
However, these evaluations often fail to consider the model’s existing knowledge, leading to unnecessary and repetitive API calls without specific prompt instructions. Furthermore, the benchmarks like AgentBoard place undue emphasis on the sequence of sub-goals, sometimes misjudging a model’s progress when redundant sub-goals are involved. This rigid focus can result in an inaccurate assessment of the model’s true capabilities in task completion.
We address all these current issues in the existing benchmark, where our focus is more on assessing the quality of the action generated by the model given the constraints in the query. Our emphasis is on understanding some critical errors of the model such as hallucinations in actions w.r.t API names or argument names, performing redundant repetitive calls, invalid format issues etc. It is important to quantify such issues to get a deeper insight into the improvements required in the model.
@@ 3 Error Patterns in Tool-Use
It’s widely understood that LLMs are prone to output errors such as hallucinations, inconsistencies, errors, etc (Wei et al., 2022; 2024; Lu et al., 2024). What is less studied are the errors exhibited in tool-use scenarios. In tool-use scenarios, typically the LLM outputs one or more function calls, including necessary arguments for the execution of the functions. For example, an LLM output for using a web search API to learn about the latest fashion can be as the following: Search[query="latest fashion", top_k=10], where Search indicates the tool and query and top_k are the context-specific arguments for executing the tool call.
@@
Note that these tool calls are particularly brittle - even a slight change to an argument, a missing argument, or incorrect tool call can produce drastically different and incorrect results downstream. Given the importance of catching these tool call errors, in this work, we characterize systematic errors generated by LLMs on tool-use tasks into seven critical error types.
@@
  • Insufficient API Calls (IAC): Unable to generate sufficient API calls, hence unable to completely fulfill the tasks provided in a query.
@@
  • Incorrect Argument Value (IAV): Generates incorrect argument values. This also includes exclusion of required arguments.
  • Incorrect Argument Name (IAN): Hallucinates argument names.
  • Incorrect Argument Type (IAT): Generates incorrect argument type.
@@
  • Repeated API Calls (RAC): Generates the exact same API Call repeatedly, causing redundant calls.
  • Incorrect Function Name (IFN) : Hallucinates function names, not part of the API list.
  • Invalid Format Error (IFE) : Generates inappropriate format.
Query Error Model Output
I need jokes sorted by score in ascending order of the category food? Can you provide me with that? IAV GPT-3.5
GT: getJokeCategory(category=‘food’, sortby= ‘asc’)
MO: getJokeCategory(category=‘workplace’, sortby = ‘asc’)
I want the genre, runtime, IMDb rating, language for ‘Endgame’? Also, include the streaming platforms available in the US. IAC GPT-4
GT: getMovie(title=‘Endgame’,type=‘movie’),                                             getMovieProvider(region = ‘US’)
MO: getMovie(title=‘Endgame’,type=‘movie’)
I need the titles and release years for horror movies with a minimum IMDb of 7 released between 1970 - 2020. IAN MLlama
GT: getMovieAdv(genre = ‘horror’, startyear = 1970, endyear = 2020, minimdb = 7)
MO: getMovieAdv(genre=‘horror’, releasestart=1970 , releaseend=2020, imdbrating =7)
I need to check if the domain ‘business.com’ is available or not IAT xLAM
GT: checkDomainAvailability(domain = ‘business.com’, availableonly=True)
MO: checkDomainAvailability(domain= ‘business.com’ , availableonly=”True”)
4 ToolScan Dataset
Figure 1: Query Generation Workflow for ToolScan. Query and API Info collected from Open Source Toolsets are given as input to an LLM with a system prompt. The LLM is asked to generate augmented queries along with the API Calls required to solve that query.
@@ 4 ToolScan Dataset
The brittle nature of LLM-generated tool calls highlights the need to quickly discover errors in LLM-generated outputs. To quickly discover tool-calling errors in LLMs, we propose a new dataset, ToolScan. It is created through a series of effective steps aimed at generating complex queries.
@@
The process begins by gathering seed queries for each API environment, sourced from benchmark tests conducted on Toolsets like ToolBench(Guo et al., 2024) and AgentBoard(Ma et al., 2024). These seed queries are then augmented using the following methods also shown in Figure 1.
@@
Constraint based Query Generation: Utilizing prompt engineering techniques, we incorporate detailed descriptions of each argument and its corresponding argument value choices to enhance the original query using GPT-4. We introduce method by adding multiple arguments and their options, making the query even more complex.
@@
Relevant Data based Query Generation: On seed queries, we instruct GPT-4 to introduce unnecessary information around the keywords while maintaining the original context of the query.
Environments Sampled Queries Avg. Turns Avg. APIs
@@
ToolScan 150 6.8 6.2
Table 3: Comparison of sampled queries, interactions, and APIs across environments.
Based on our required criteria of creating complicated queries we generate 600 queries. We filter these queries based on feasibility evaluation, constraint validation and diversity. For example we remove infeasible queries such as incorrect function name, incorrect arguments, etc which leaves about 150 queries. More details in Appendix A.1.4.
Table 3 describes the distribution of queries across different environment APIs. We also list down the average number of interactions along with the queries sampled per environment after augmentation.
5 Framework
@@
As shown in Figure 2, our benchmark follows a deterministic framework where the agent will begin with Instruction $i$, Query $q$, and a sequence of API-list $A$ (resembling to the action space), provided to successfully solve the provided query. The agent then based on the policy of the language model will perform an action $a_{t}$ in the current $s_{t}$, leading to a determined state transition $s_{t+1}$.
A constructive feedback mechanism f is incorporated to help assist the agent for the following issues:
  • Verifies the generated action is correctly parsed and adheres to the format instructions provided in the prompt.
  • Determines if the generated action is in the specified action space. If not, enumerates the list of available actions that can be used to resolve the query.
  • Assesses whether the generated action is invoked with the appropriate arguments. If not, provide a detailed list of the available arguments for the chosen action, along with their respective descriptions.
  • Ensures that the generated action is called with arguments of the correct type. If not, specify the correct argument type required for proper execution of the function.
@@
Figure 2: Systemic workflow of an agent interacting with the environment assessing the prior information and the predefined goal to determine its next action.
@@ 6 Metrics
Let $Q$ be the set of queries. For each query $q_{i}\in Q$, let $N$ represents the maximum steps the model is permitted to process the query, while $G_{i}$ denotes the Ground Truth labels for the $i^{th}$ query $q_{i}\in Q$. Metric definitions based on above are as follows:
@@
  • For error patterns IFN, IAN, IAT, IFE and RAC we calculate the percentage of API Calls executed without respective error as $\sum_{i=1}^{|Q|}1-\frac{e^{i}}{N}$, where $e^{i}$ is the number of API calls with the error pattern in $i^{th}$ query $q_{i}$.
  • For error patterns IAC and IAV, let $\{g\in G_{i}\}$ be the set of possible trajectories for query $q_{i}$. Then we define our metric as for each error pattern as
    $\sum_{i=1}^{|Q|}1-\frac{\min_{g\in G_{i}}\mathrm{e^{i}_{g}}}{N}$
    where $e_{g}^{i}$ are the number of API calls with error pattern when evaluated in comparison to ground truth trajectory $g\in G_{i}$ for query $q_{i}\in Q$.
  • For each query, we calculate success rate by checking if the model is able to successfully complete the query.
@@
It is crucial to emphasize that while the metrics indicate error patterns, their calculations are designed to be directly proportional to the success rate. Higher scores on the error metrics represent fewer issues generated by a model in that respective error pattern.
7 Experiments
We’ve conducted a comprehensive assessment of several performant LLMs, including GPT-4 (Achiam et al., 2023), GPT-3.5-Turbo from OpenAI, Meta-Llama3-8b (Dubey et al., 2024), DeepSeek-R1-Distill Models (Guo et al., 2025), Code-Llama-13B (Rozière et al., 2024), Vicuna-13B-16k (Chiang et al., 2023), Mistral and Mixtral models from Mistral AI (Jiang et al., 2023) (Jiang et al., 2024), and xLAM (Zhang et al., 2024). Interestingly, while these models all perform well on ToolBench(Guo et al., 2024), their failure patterns are different. Details of the error patterns discovered using ToolScan are shown in Table 2.
We incorporate a one-shot in-context example along with task-specific instructions in our prompt setup. For each open-weight model, we evaluate the chat-optimized version when available, unless otherwise indicated. For models without a dedicated chat version, we apply standardized prompt templates tailored for publicly accessible versions. We ensure that each prompt includes precise format instructions to enable the model to generate API calls in a structured and parsable format. To account for server inconsistencies, we execute multiple rounds per API call, thus minimizing potential disruptions and ensuring consistent evaluation outcomes.
8 Results
From our results in Table 4, we observe that LLMs specifically fine-tuned for API calling are generally less likely to produce false information, often known as ”hallucination” when dealing with function names and argument names. On the contrary, models like Vicuna (Chiang et al., 2023) (Zheng et al., 2024) that are designed for chat-based interactions have a higher rate of inaccuracies, particularly incorrect argument names and function names. It suggests that the tuning process, specifically for API calls, can lead to a substantial reduction in hallucination.
We observe that Insufficient API Calls (IAC) is the most common error pattern. Most models have a significantly low IAC score. We observed two reasons for it: (1) Models like GPT only solves part of the query. If there are multiple questions asked in the same query, GPT ends after answering the fist query. (2) Models are unable to distinguish similar APIs and ignore the given constraints. For example, model will pick get_patent_with_title instead of get_patent_with_title_and_date and hence only return the patents with title and not date.
We also observe that smaller models which are not tuned for function calling such as Meta-Llama3-8b, DeepSeek-R1-Distill-Qwen-7B etc. perform Repeated API Calls (RAC) despite clear instructions to avoid repetition. They get stuck in a loop based on their internal belief of solving the query and call the same API multiple times even if they are failing. We also see a co-relation between RAC and IAC between these models which suggests that RAC often leads to IAC.
For Incorrect Argument Value (IAV), we observe that models tend to ignore optional arguments like sorting, limit etc and tends to go with the default values set in the functions. This can be seen in models like Code-Llama-13B, Mixtral-8x22B-Instruct-v0.1 and DeepSeek-R1-Distill-Qwen-14B. This error is especially prevalent in smaller models like Mistral 7B because it ignores ”required” arguments explicitly mentioned in the APIList.
Models like Mixtral-8x22B-Instruct-v0.1 also face errors in specifying correct argument type, confusing string for int and vice-versa. This increase the Incorrect Argument Type (IAT) error and hence reducing IAT score. Depending on the type of training done on models they tend to exhibit different outcomes in error patterns. For instance, DeepSeek-R1 distill models are specifically trained on MATH tasks with profound focus on improving their reasoning. Hence, on tool-use tasks these models face difficulty in following format instructions (low IFE) and spend significant chunk of tokens in the thinking process.
Most striking is the superior performance of GPT-4, achieving the highest success rate among all models evaluated. This could potentially suggest that larger models harnessing the power of the scaling law and quality pre-training in the base model remain critical to the success of agent applications. Another noteworthy observation is the exemplary performance of the Code-Llama-13B model (Rozière et al., 2024), surpassing many other generically-purposed models. Our hypothesis is that there may be some inherent similarity between coding-associated tasks and function calls. And that could be beneficial for models designed for coding purposes, thereby enhancing their performance on function-calling tasks.